# Connecting the Dots: Evaluating Abstract Reasoning Capabilities of LLMs Using the New York Times Connections Word Game

### v3

@@ 1 Introduction
Abstract reasoning represents a person’s ability to solve problems, identify patterns, and work with logical systems Barrett et al. (2018); Johnson et al. (2021). While the performance of large language models (LLMs) on arithmetic and language-based commonsense reasoning benchmarks has been the subject of recent analyses, it is unclear whether these LLMs possess abstract reasoning capabilities that are often challenging even for humans Xu et al. (2023). Given its nature, we choose the NYT Connections Game as a test bed for investigating the abstract reasoning capabilities of both humans and large language models (LLMs). We collect 200 distinct Connection games and test the capabilities of four state-of-the-art large language models, namely Google’s Gemini 1.5 Pro Team et al. (2023), Anthropic’s Claude 3 Opus Anthropic (2024), OpenAI’s GPT4-Omni OpenAI (2023), and Meta’s Llama 3 70B AI@Meta (2024) and compare them with human performance.
@@ 2 Related Work
Advancements in LLMs have led to a growing interest in exploring their potential to take on intricate and conceptual challenges by generating linguistic sequences. These models have shown potential in gaming by serving as players Wang et al. (2023a); Tsai et al. (2023); Ciolino et al. (2020); Bakhtin et al. (2022); Noever et al. (2020), non-player characters (NPCs) Park et al. (2023); Urbanek et al. (2019), and generating game content Todd et al. (2023); Wang et al. (2023b); Sudhakaran et al. (2024); Ammanabrolu and Riedl (2021).
Recent research has explored applying large language models (LLMs) and other natural language processing (NLP) techniques to solve and generate text-based puzzles. Zhao and Anderson (2023) use LLMs to tackle and create the weekly Sunday Puzzles featured on National Public Radio (NPR). When presented with multiple-choice questions, GPT-3.5 attained an accuracy of up to 50%. However, when asked to generate novel and engaging puzzles, the model encounters challenges. Compared to the Connections game, NPR’s weekly puzzles tend to emphasize character-level word transformations and relatively common references, relying less on encyclopedic, associative, or semantic knowledge. Rozner et al. (2021) examined the potential of using "cryptic crossword" clues as an NLP benchmark. Wallace et al. (2022) propose automatic ways of solving crossword puzzles by generating answer candidates for each crossword clue using neural question answering models and combining loopy belief propagation with local search to find full puzzle solutions. Our work builds upon these efforts by utilizing the NYT Connections puzzle as a means to investigate the abstract reasoning capabilities of state-of-the-art LLMs.
@@ 4.2 Humans as Game Players
We designed a human evaluation interface and randomly sampled 100 games from our test set. Appendix E has more information about the interface. The first screen displays a shortened version of the instructions from the LLM’s prompt so as to not overwhelm the human players. To ensure that the humans solve the game in a manner comparable to the LLMs setup, they were given one try to fully solve the game (i.e., make all 4 categorizations).
@@ 5.1 LLM performance
Overall, we find that GPT-4o performs best across all 200 games. Figure  2 shows the unweighted clustering scores for all 4 LLMs. GPT-4o has the lowest percentage of games in which it made no correct clustering (53 out of 200) and the most games solved perfectly (16 out of 200). Claude 3 Opus is a close second for perfectly solved games at 13 out of 200. Gemini 1.5 Pro performs the worst overall. While it could not make any correct clusters for half of the games, it was able to solve 9 games perfectly, outperforming Llama 3 70B’s 6 games solved perfectly.

### v4

@@ 1 Introduction
Abstract reasoning represents a person’s ability to solve problems, identify patterns, and work with logical systems Barrett et al. (2018); Johnson et al. (2021); Ji et al. (2022). While the performance of large language models (LLMs) on arithmetic and language-based commonsense reasoning benchmarks has been the subject of recent analyses, it is unclear whether these LLMs possess abstract reasoning capabilities that are often challenging even for humans Xu et al. (2023). Given its nature, we choose the NYT Connections Game as a test bed for investigating the abstract reasoning capabilities of both humans and large language models (LLMs). We collect 200 distinct Connection games and test the capabilities of four state-of-the-art large language models, namely Google’s Gemini 1.5 Pro Team et al. (2023), Anthropic’s Claude 3 Opus Anthropic (2024), OpenAI’s GPT4-Omni OpenAI (2023), and Meta’s Llama 3 70B AI@Meta (2024) and compare them with human performance.
@@ 2 Related Work
The growing popularity in Large Language Models have led to an exciting array of research using natural language processing techniques for text-based games. Recent work has studied whether these models can act as players in agentic environment Huang et al. (2024); Wang et al. (2023); Wu et al. (2023); Noever et al. (2020) or conversational setting Qiao et al. (2023). In addition to acting as players researchers have also tested the abilities of transformer based language models in generating games Ammanabrolu and Riedl (2021); Todd et al. (2023); Sudhakaran et al. (2024); Hu et al. (2024); Chen et al. (2023)
Recent research has explored applying large language models (LLMs) and other natural language processing (NLP) techniques to solve and generate text-based puzzles. Wallace et al. (2022) propose automatic ways of solving crossword puzzles by generating answer candidates for each crossword clue using neural question answering models and combining loopy belief propagation with local search to find full puzzle solutions.Zhao and Anderson (2023) use LLMs to tackle and create the weekly Sunday Puzzles featured on NPR. They release PUZZLEQA, a multiple-choice dataset comprising 15 years of on-air puzzles and show that while ChatGPT can solve these questions with an accuracy of around 50% they still struggle at generating novel and engaging puzzles. Unlike Connections game, PUZZLEQA relies on character-level word transformations compared to encyclopedic, associative, or semantic knowledge. Rozner et al. (2021) examined the potential of using "cryptic crossword" clues as an NLP benchmark.
Contemporaneous work from Todd et al. (2024) test OpenAI’s GPT3.5 and GPT4’s abilities in solving the Connection puzzles using Chain-Of-Thought prompting. Our work is similar to theirs in a way that it utilizes the NYT Connections puzzle as a means to investigate the abstract reasoning capabilities of state-of-the-art LLMs. However, in addition to GPT4 we also test 3 other state of the art LLMs (both open and closed weight) as well as humans(both novice and experts). Our contribution not only focuses on the abilities of LLMs to solve the games but we also study the knowledge types required to solve the Connections game.
@@ 4.2 Humans as Game Players
We designed a human evaluation interface and randomly sampled 100 games from our test set. Appendix E has more information about the interface. The first screen displays a shortened version of the instructions from the LLM’s prompt so as to not overwhelm the human players. To ensure that the humans solve the game in a manner comparable to the LLMs setup, they were given one try to fully solve the game (i.e., make all 4 categorizations). This aligns with Todd et al. (2024)’s challenge mode.
@@ 5.1 LLM performance
Overall, we find that GPT-4o performs best across all 200 games. Figure  2 shows the unweighted clustering scores for all 4 LLMs. GPT-4o has the lowest percentage of games in which it made no correct clustering (53 out of 200) and the most games solved perfectly (16 out of 200). Our results with GPT4o is similar to those reported by Todd et al. (2024) (12.5% success rate by GPT4-turbo). Claude 3 Opus is a close second for perfectly solved games at 13 out of 200. Gemini 1.5 Pro performs the worst overall. While it could not make any correct clusters for half of the games, it was able to solve 9 games perfectly, outperforming Llama 3 70B’s 6 games solved perfectly.